OpenStudy (samsan9):

what are the x-intercepts of the graph y=3(x+1)(x-6)?

OpenStudy (samsan9):

A.) (0,1) and (0,-6) B.) (1,0) and (-6,0) C.) (-1,0) and (6,0) D.) (0,-1) and (0,6)

OpenStudy (anonymous):

y=3(x+1)(x-6) when the line intercepts x, y=0 therefore 3(x+1)(x-6)=0 expand the eqn, we have 3x^2-15x-18=0 we have x=6 and x=-1

OpenStudy (anonymous):

expand?

OpenStudy (samsan9):

do we plug in the answer choices in to the equation to see if they equal each other?

OpenStudy (anonymous):

yeah that's how you check to see if your answer is correct.

OpenStudy (anonymous):

\[3(x+1)(x-6)=0\iff x+1=0,x-6=0\]

OpenStudy (samsan9):

so what would you the answer could be so i i just plugin -1and 6?

OpenStudy (anonymous):

the answer would be 0 if you plug in the right ans

OpenStudy (samsan9):

so it has to be equal to 0for it to be correct?

OpenStudy (anonymous):

yah

OpenStudy (samsan9):

thanks :D